OpenStudy (anonymous):

Solve the system of equations: what is the x-value for the solution? 2x + 7y = 3 and x = 1 - 4y A) -1 B)5 C)4 D)3

OpenStudy (johnweldon1993):

you want the x-value...so I would solve 1 of these equations for y...then plug it into the other equation 2x + 7y = 3 7y = 3 - 2x y = 3/7 - 2/7x this is y....so plug it into the second equation x = 1 - 4y replace "y" with what you GOT for y x = 1 - 4(-2/7x + 3/7) x = 1 + 8/7x - 12/7 get the 2 x's together x - 8/7x = 1 - 12/7 -1/7x = -5/7 multiply both sides by -7 x = 35/7 x = 5 If i am not correct...anyone please correct me :)

OpenStudy (anonymous):

2x + 7y=3, x = 1-4y 2(1-4y)+7y=3 2-8y+7y=3 2-1y=3 2-y=3 y=3/-2 y=-1

OpenStudy (anonymous):

Thank you!

OpenStudy (johnweldon1993):

@Tanieshia right....if x = 5 like I said...then the y would = -1 you wanted the x number @ally44 correct? so that would be the 5 like I said

OpenStudy (anonymous):

ok @ johnweldon, i got u since they already gave x i thought we were looking for y

OpenStudy (johnweldon1993):

lol no problem...now she would see how to solve for y as well :)

OpenStudy (anonymous):

it's B

OpenStudy (anonymous):

lol ok so she needed x and not y? ok my bad

OpenStudy (anonymous):



OpenStudy (anonymous):

johnweldon1993 can i have your views on this one pls

OpenStudy (johnweldon1993):

hmmm well the lines of symmetry would be you can fold it over and have the SAME shape... so you can fold it in half...thats 1 you can fold it in half down...thats 2 you can fold it into a triangle shape 2 different ways...thats 4 now hmm...anyone else add any? lol

OpenStudy (anonymous):

thank you so much